. An administrator of a popular web site is told that a new server can handle 41,000 "hits" (users accessing the site) per second. The web site currently experiences a peak demand of about 105,000 hits per second; but every month, the peak demand increases by 2800 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 24 months.

Answer by gonzo(654)

assuming the new server is already taking care of the current demand, then 3 servers must already be in service with a capacity of 3 * 41000 = 123000 hits.
-----
if you are currently experiencing 105000 hits per second, then you have spare capacity of 123000 = 105000 = 18000 hits per second.
-----
your expected growth in the next 24 months is 24 * 2800 = 67200 hits per second.
-----
subtract your spare capacity from the expected growth and you have 67200 - 18000 = 49200 growth that you don't have capacity for.
-----
divide that growth by 41000 and you will need 1.2 additional servers.
-----
since you can't buy .2 of a server, you will need 2 additional servers.
-----
not sure if this answers your question.
if it doesn't, then resubmit your question to get other answers.
-----